Mikaela1509:

7n + 8 = 71 Help please

Joe348:

so subtract 8

Rylee88:

7(What) + 8 = 71? Find the variable.

Joe348:

rlyee stop I got this

Joe348:

@joe348 wrote:
so subtract 8
rn=63

Rylee88:

If I can help I will =)

IrispurpleClub:

she is find

Joe348:

@joe348 wrote:
@joe348 wrote:
so subtract 8
rn=63
then you want to divide both sides by 7

Joe348:

@joe348 wrote:
@joe348 wrote:
@joe348 wrote:
so subtract 8
rn=63
then you want to divide both sides by 7
63 divided by 7 equals 9

IrispurpleClub:

math is so hard

Mikaela1509:

ok this is still very confussing

Rylee88:

7 𝑛 + 8 = 7 1 7n+8=71 7n+8=71 Solve 1 Subtract 8 8 8 from both sides of the equation 7 𝑛 + 8 = 7 1 7n+8=71 7n+8=71 7 𝑛 + 8 − 8 = 7 1 − 8 7n+8{-8}}=71 7n+8−8=71−8 2 Simplify Subtract the numbers Subtract the numbers 7 𝑛 = 6 3 7n=63 7n=63 3 Divide both sides of the equation by the same term 7 𝑛 = 6 3 7n=63 7n=63 7 𝑛 7 = 6 3 7=763​ 4 Simplify Cancel terms that are in both the numerator and denominator Divide the numbers 𝑛 = 9 n=9 n=9 Solution 𝑛 = 9

Joe348:

i got you if its still confusing

Joe348:

so you want to combine like terms

Joe348:

which is 8 and 71

Joe348:

\[7n+8=71\] \[-8 on both sides\] \[7n=63\] divide both sides by 7 since there is no varaibles left \[63\div 7 = 9\]

Mikaela1509:

Oh! thank you so much

Mikaela1509:

I really appreciate it